Fermi National Laboratory
 

Accelerator Update for March 9, 2007 - March 12, 2007

 

Friday March 9

The day shift began with Operations monitoring the Recycler stash (319.1E10), with the Antiproton Source (Pbar) stacking (44.9E10), with the Booster, Linac, Main Injector (MI), and Recycler operating normally, with MiniBooNE off, with NuMI and the Meson MTest experiment taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) waiting the colliding beams.

At 8:23 AM, cryo system technicians reported a problem with the Pbar dry engine. They began repairs.

At 8:41 AM, Operations reported that the Linac Buncher had tripped off. Experts investigated.

Operations put the Tevatron (TeV) into shot setup around 9 AM.

Recycler experts began a study period at 10:55 AM.

Operations established store 5268 at 11:08 AM with an initial luminosity of 267.13E30.

A Pbar expert began a flying wire study at 11:13 AM.

At 12:26 PM, a Booster RF station (BRF10) tripped off. It reset.

At 12:49 PM, Operations reported that the MI BLMs were aborting the MI beam due to bad beam positions. The NuMI beam permit was also tripping due to this. MIRF14 had tripped off earlier and this might have contributed to these problems. MI experts were investigating. Operations lowered the beam intensity.

At 1:06 PM, RF experts reported that they wanted to work on BRF7. Operators bypassed the station and brought BRF19 online to compensate.

Recycler experts reported that they had completed their studies at 2:50 PM.

MI experts began a MI-8 collimator study at 3:26 PM. This study reduced beam intensity.

Operations reported at 3:47 PM that the Recycler had stashed antiprotons once during the day shift.

Around 4 PM, the NuMI beam permit went down due to a problem with its target BLM gas flow. A NuMI expert was called in to investigate.

A NuMI expert reported 5:37 PM that she had fixed the target BLM gas flow problem.

At 6:10 PM, a Tevatron Electron Lens (TEL2) tripped off due to a vacuum interlock trip. It reset okay.

The MTest experiment reported at 6:29 PM that they had finished using the beam for the night.

A Pbar expert began an AP2 orbit study.

At 9:10 PM, BRF6 tripped off. It reset.

At 9:54 PM, LRF3 tripped off for the first of two times in nine minutes. It reset okay.

Operations reported at midnight that the Recycler had stashed three times during the evening shift.

Saturday March 10

The midnight shift began with Operations monitoring store 5268 (71.07E30) and the stash (173.2E10), with Pbar stacking (38.1E10), and with NuMI taking beam.

Operations established beam to MTest at 6:21 AM.

Operations terminated store 5268 at 8 AM.

Operations reported at 8:21 AM that the Recycler had stashed four times during the midnight shift.

Operations put the TeV into shot setup at 8:22 AM.

Operations established store 5269 at 10:14 AM with an initial luminosity of 220.4E10.

At 10:48 AM, Operations reported that a Linac quadrupole power supply’s (L:QPS206) output current was noisy. Operators swapped control cards.

At 11:35 AM, Operations reported that a Pbar Debuncher kicker (D:IKIK) began to fail intermittently. By 11:46 AM, it resume normal running.

At 12:15 PM, Operations reported that the swap of control cards for L:QPS206 didn’t resolve anything. Operators started swapping other control cards, which did resolve the problem.

At 12:35 PM, Operations reported a fire alarm for the crass-gallery and the MCR. Operations halted beam and evacuated the control room. The Fire Department allowed the operators to return to the MCR at 12:50 PM. Normal beam operations resumed.

At 5:38 PM, the Linac Buncher tripped off. Operators had the Buncher back up within four minutes.

MTest reported at 6:36 PM that they were done with the beam for the night.

Operations reported at 7:48 PM that the Recycler had stashed four times during the last 12 hours.

Operations reported the following beam status at 8 PM: Store 5269 had a luminosity of 74.16E30, the stash was at 170.6E10, and the stack was at 23.3E10.

At 9:15 PM, Operations reported that D:IKIK went erratic for about five minutes.

At 11:59 PM, a Pbar power supply (M:HT105) pulled the Pbar beam permit. Operations did a restore on a Pbar tuning program and this resolved the problem.

Sunday March 11

The midnight shift began with Operations monitoring store 5269 and the stash, with Pbar stacking, and with NuMI taking beam.

At 3:26 AM, M:HT105 pulled the Pbar beam permit for the first of five times in about three hours. It reset okay each time. Operations thought that the tuning program was driving this power supply too hard. Operators investigated.

At 6:56 AM, a number of Proton Source, MI, and MuCool devices tripped off. The source of the problem was a Booster CAMAC crate that wasn’t responding. This crate held interlock cards for all these areas. Its power supply had failed. Operators replaced the power supply. Beam resumed at 7:19 AM.

At 7:27 AM, Operations reported that beam was not making it to the Pbar Debuncher. Operators investigated.

Operations established beam to MTest at 7:47 AM.

Operations terminated store 5269 at 8 AM.

At 8:07 AM, Operators reported that the Pbar lithium lens’ voltage was down. Operators turned it on and off a few times and the lens resumed normal operations.

Operations reported at 8:15 AM that the Recycler had stashed three times during the last 12 hours.

Operations put the TeV into shot setup at 8:26 AM.

At 8:40 AM, a Linac Klystron RF station (KRF5) tripped off for the first of five times in 27 minutes. It reset each time.

At 9:30 AM, Operations reported that the Pelletron looked like it had discharged, but they weren’t sure. Operations contacted an expert. Operations reported at 10:06 AM that this was a communication problem and that with the help of an expert they resolved it. The Pelletron was on and working normally by 10:45 AM. Operations resumed shot setup.

At 12:31 PM, shortly after all the antiprotons were loaded, the TeV suffered a sector E48 quench at the start of ramping. Operations contacted an expert.

TeV experts reported at 2:23 PM that there was no direct evidence of what caused the E48 quench, but they suspect the sextupoles were involved.

Operations loaded a 4x0 wet squeeze into the TeV at 4:28 PM.

Operations put the TeV into shot setup at 5:49 PM.

At 7:20 PM, Operations reported at that the LCW temperature in MI-20 was too high. Operators thought it was a heat exchanger problem.

At 7:46 PM, Operations reported that the MI-20 problem was due to a clogged strainer or heat exchanger. Operators reduced the beam intensity for the area and then let the Duty Mechanic investigate further.

Operations reported at 8:08 PM that the Recycler had stashed two times during the last 12 hours.

MTest reported at 9:20 PM that they were done with the beam for the night.

Operations reported 9:31 PM that the Duty Mechanic switched the MI-20 pond pump to high speed and brought the temperatures down.

Operations established store 5274 at 9:36 PM with an initial luminosity of 153.49E30.

Monday March 12

The midnight shift began with Operations monitoring store 5274 and the stash, with Pbar stacking, and with NuMI taking beam.

At 12:49 AM, a TeV RF station (TRF4) tripped off and caused a drop in luminosity.

At 1:26 AM, KRF5 tripped off. It reset.

At 2:02 AM, BRF6 tripped off. It reset.

At 2:20 AM, BRF16 tripped off. It reset.

At 4:34 AM, LRF3 tripped off. It reset.

At 5:57 AM, LRF1 tripped off. It reset.

Operations made Pbar ready for access at 7:54 AM.

Operations reported at 8:11 AM that the Recycler had stashed five times during the last 12 hours.

The Plans for Friday

The plans for today are to allow about four hours of maintenance work in the TeV and Pbar. NuMI will continue to receive beam. The next store will be established on the evening shift.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/12/2007    email Fermilab